# On canonical differential equations for Calabi-Yau multi-scale Feynman integrals

### v1

@@ 1 Introduction
Feynman integrals are central quantities in Quantum Field Theory, as they are essential building blocks for computing scattering amplitudes which are one of the bridges between theory and experiment Travaglini:2022uwo . Integrals of similar type have also appeared in the description of the scattering of compact objects in classical General Relativity in the so-called post-Minkowskian expansion Goldberger:2004jt ; Porto:2016pyg ; Mogull:2020sak . It is therefore necessary to develop sophisticated techniques which help us compute this type of integrals, in particular, by understanding their mathematical structure and underlying geometrical information. It turns out that at the core of all these integrals are classes of special functions related to complex manifolds where the simplest is the Riemann sphere where multiple polylogarithms (MPLs) Kummer ; Remiddi:1999ew ; Goncharov:1995 ; Goncharov:1998kja ; Vollinga:2004sn ; Duhr:2011zq appear. However, diagrams needed for cutting-edge computations give rise to special functions beyond the ones that live on the Riemann sphere. These functions beyond MPLs are associated to periods of different geometries. The first example being the massive sunrise diagram in even dimensions which can be associated to an elliptic curve (which is isomorphic to a torus) Bloch:2013tra ; Broadhurst:1987ei ; Bauberger:1994by ; Bauberger:1994hx ; Laporta:2004rb . Recent progress has been made to go beyond the torus, on the one hand by studying Feynman integrals associated to higher genus Riemann surfaces Georgoudis:2015hca ; Duhr:2024uid and their function space DHoker:2023vax ; Baune:2024biq ; DHoker:2025szl . On the other hand by studying Feynman integrals associated to higher dimensional Calabi-Yau manifolds Primo:2017ipr ; Klemm:2019dbm ; Bourjaily:2018yfy and their function space Duhr:2025ppd ; Duhr:2025tdf .
@@
In recent years, substantial effort has been dedicated to extend the definition of canonical differential equations to Feynman integrals beyond polylogarithms, and much progress has been made, in particular in the elliptic and one-parameter Calabi-Yau cases Gorges:2023zgv ; Pogel:2022yat ; Adams:2016xah ; Adams:2018bsn ; Pogel:2022ken ; Frellesvig:2021hkr ; Dlapa:2022wdu ; Frellesvig:2023iwr ; Chen:2025hzq ; Pogel:2022vat . In particular, ref. Pogel:2022vat provides an ansatz to bring equal mass banana integrals, with underlying single-scale Calabi-Yau geometry, at any loop order, into $\epsilon$-factorised form. Moreover, very recently in ref. Duhr:2025lbz , a generalisation of the method of ref. Gorges:2023zgv has been developed for generic integrals with an underlying Calabi-Yau geometry (and beyond), providing explicit calculations for cutting edge Calabi-Yau one-scale integrals. The applicability of this algorithm has been explicitly demonstrated for many state-of-the-art problems containing integrals of elliptic type Gorges:2023zgv ; Becchetti:2025rrz ; Becchetti:2025oyb ; Duhr:2024bzt ; Forner:2024ojj ; Marzucca:2025eak and for their generalisation to Calabi-Yau varieties Duhr:2024bzt ; Forner:2024ojj ; Klemm:2024wtd ; Driesse:2024feo ; Driesse:2024xad ; Duhr:2025lbz and higher-genus surfaces Duhr:2024uid .
In this paper, we propose a generalisation of the method introduced in ref. Pogel:2022vat , to multi-scale Calabi-Yau integrals. In the preparation of this work, we have realised that our generalisation coincides with the method from refs. Gorges:2023zgv ; Duhr:2025lbz , applied to multi-scale Calabi-Yau Feynman integrals. In fact, this is the first application of refs. Gorges:2023zgv ; Duhr:2025lbz to this class of integrals. To support our claim, we derive locally at the MUM-point, for the first time, a canonical differential equation for banana integrals with unequal mass.
This paper is organised as follows. In section 2 we set our conventions and review the instruments that will be needed for the procedure we propose. Following this, in section 3, we present our procedure for multi-variate Feynman integrals with underlying Calabi-Yau geometry. In section 4, we recompute a canonical form for the unequal mass sunrise integral using the method presented before. Finally, in section 5, we apply our construction to compute a canonical form for integrals with underlying multi-variate Calabi-Yau geometry, namely, the three-loop banana with two unequal masses (two-parameters K3) and the four-loop banana with two unequal masses (two-parameters CY 3-fold). Additionally, we include two appendices in which we prove the validity of our procedure for the elliptic and Calabi-Yau two-fold case. The main quantitative results will be compiled in the ancillary files sunrise.nb, banana_3_1.nb, banana_2_2.nb and banana_4_1.nb.
@@ 2.1 Feynman integrals and their differential equations
where $\alpha_{jk}\beta_{jk}$ can be chosen to lie in $\{0,\pm 1\}$. Any integral of the form (1) can be reduced to a finite linear combination of master integrals (MIs) which has fixed $\nu_{j}$. Every integral belongs to a sector, defined as the set of integrals from the family that share exactly the same propagators, i.e., $I_{\nu_{1},\dots,\nu_{r}}$ and $I_{\nu^{\prime}_{1},\dots,\nu^{\prime}_{r}}$ belong to the same sector if $\theta(\nu_{i})=\theta(\nu^{\prime}_{i})$, for all $1\leq i\leq r$, where $\theta$ denotes the Heaviside step function, $\theta(x)=1$ if $x>0$ and $\theta(x)=0$ otherwise. Determining these MIs and their associated differential equation can be achieved through the Laporta algorithm, where different software implementations are available Lee:2012cn ; Smirnov:2008iw ; Smirnov:2019qkx ; Klappert:2020nbg ; Maierhofer:2017gsa . This leads to a system of linear first-order differential equations for the vector of MIs $\bf I$ that can be shown to be Kotikov:1991hm ; Kotikov:1991pm ; Henn:2013pwa ; Kotikov:1990kg ; Gehrmann:1999as :
@@
${\bf{M}}\left({\bf z};\epsilon\right)$ $=\mathcal{P}\exp\left(\epsilon\int_{{\bf z}_{0}}^{{\bf z}}\tilde{\bf A}({\bf z}^{\prime})\right){\bf{M}}({\bf z}_{0};\epsilon)\,,$ (6)
@@ 2.2 Feynman integrals and their geometry
In all the examples we consider, we find that the maximal cut only features one geometry. In cases where there are more than one geometry, we expect the differential equation to split into different sub-blocks, provided that we choose a good starting basis as predicted by the MHS. Equivalently as putting the propagators on-shell, a maximal cut of a Feynman integral is obtained by evaluating the integrand in Baikov representation on a contour that encircles all the propagator poles. The integration contour varies according to the geometry of the Feynman integral considered. For instance, at one loop we encounter complex curves of genus zero, namely Riemann spheres. In this case, maximal cuts, in the appropriate integer dimension, evaluate to algebraic functions and the differential one-forms obtained in the differential equation are the so-called $dlog$-forms, which after integration evaluate to multiple polylogarithms (MPLs) Kummer ; Remiddi:1999ew ; Goncharov:1995 ; Goncharov:1998kja ; Vollinga:2004sn ; Duhr:2011zq .
@@
where $f_{k}(\tau)$ is a modular form of weight $k$ for some congruence subgroup and $g^{(k)}(z,\tau)$ are the coefficients of the expansion of the Kronecker function. A similar analysis applies when the degree of the polynomial under the square root in the maximal cut is three Broedel:2017kkb ; Broedel:2017siw .
@@ 2.3 Picard-Fuchs operators
[…]
@@
which generalises the $\tau$ parameter of an elliptic curve and gives a local (at the MUM-point), canonical variable on the moduli space of Calabi-Yau varieties. The dots denote local analytic terms in $z$. The inverse of this map is also known as the mirror map
@@
$N_{0}=1\,,\quad N_{j+1}=x\dfrac{\partial}{\partial x}\dfrac{1}{N_{j}(\psi_{j}^{(l)})}N_{j}\,,$ (23)
@@ 2.4 Hodge filtration & Griffiths transversality
can be evaluated in terms of complete elliptic integral of first kind $K(k)$ and $K(1-k)$ for $1\leq z\leq\infty$ through:
@@
Introducing $\tau=\Pi_{12}(z)/\Pi_{11}(z)$ and using it as the complex modulus, we can express the periods of the elliptic curve as modular forms of weight one in $\tau$.
@@
[…]
@@
$\mathbf{\Pi}(\mathbf{z})^{T}\mathbf{\Sigma}\partial_{\bf z}^{\bf k}\mathbf{\Pi}(\mathbf{z})=\int_{M_{l}}\Omega\,\wedge\,\partial_{\bf z}^{\bf k}\Omega(\mathbf{z})=\begin{cases}0&\text{for }0\leq r<l,\\ C_{\mathbf{k}}(\mathbf{z})&\text{for }|k|=l,\end{cases}$ (42)
@@
which implies that we can find $22\hat{h}+$ 2 independent periods
@@
where the subscript determines the logarithmic power of the period in the Frobenius basis. These periods have been chosen by requiring the intersection matrix $\Sigma$ to be antidiagonal
$\mathbf{\Sigma}=\begin{pmatrix}0&\mathds{1}_{\hat{h}\times\hat{h}}\\ -\mathds{1}_{\hat{h}\times\hat{h}}&0\end{pmatrix}\,.$ (47)
@@ 2.5 Abelian differentials of first, second and third kind
where we pick the root ordering $1<z<\infty$. Topologically this corresponds to two Riemann spheres connected by two branch cuts, which we can choose to go from 00 to $1$ and from $z$ to $\infty$. If we allow differentials of the third kind, with non-vanishing residue at a marked point $c$, this corresponds to a puncture on each of the Riemann spheres at the marked point $c$. Explicitly one gets the following forms of first, second and third kind:
$\left\{\begin{array}[]{ll}\dfrac{\text{d}x}{y}&\text{: differential of the first kind},\vspace{2mm}\\ \dfrac{x\text{d}x}{y}&\text{: differential of the second kind},\vspace{2mm}\\ \dfrac{\text{d}x}{y(x-c)}&\text{: differential of the third kind,}\end{array}\right.$ (61)
@@ 2.6 $\epsilon$-factorising L-loop banana integrals of equal mass
[…]
@@ 3 Algorithm to construct a canonical basis
First of all, we need to identify the geometry of the block we are looking at. This can either be done with a Baikov analysis, or finding the periods of the Picard-Fuchs operator. For cases like the banana integrals in two dimensions, the corner integral ($I_{1,\dots,1}$) evaluates to periods of a Calabi-Yau manifold on its maximal cut in $D=2$ dimensions. This is, however, not always the case and one may have to shift dimensions or take linear combinations of the available master integrals to make the underlying geometry manifest.
@@
[…]
@@
The basis we propose has, for the examples we computed, given a canonical form which displays: [^9] So far there is no clear consensus on a definition of canonical differential equation beyond MPLs. In ref. Duhr:2025lbz , a definition was proposed, which includes the independence of the forms in addition to our defining properties. Checking the independence of the forms goes beyond the scope of this paper.
@@
[…]
@@
and we additionally require the differential equation for $\mathbf{I}_{\mathbf{B}}$ to take the form in Eq. (81). Now let us concretely show how this works for $l=1,2,3$ and then propose a generalisation for ${l>3}$.

### v2

@@ 1 Introduction
Feynman integrals are central quantities in Quantum Field Theory, as they are essential building blocks for computing scattering amplitudes which are one of the bridges between theory and experiment Travaglini:2022uwo. Integrals of similar type have also appeared in the description of the scattering of compact objects in classical General Relativity in the so-called post-Minkowskian expansion Goldberger:2004jt; Porto:2016pyg; Kalin:2020mvi; Mogull:2020sak; Kalin:2020fhe; Kalin:2022hph; Jakobsen:2022psy; Dlapa:2024cje. It is therefore necessary to develop sophisticated techniques which help us compute this type of integrals, in particular, by understanding their mathematical structure and underlying geometrical information. It turns out that at the core of all these integrals are classes of special functions related to complex manifolds where the simplest is the Riemann sphere where multiple polylogarithms (MPLs) Kummer; Remiddi:1999ew; Goncharov:1995; Goncharov:1998kja; Vollinga:2004sn; Duhr:2011zq appear. However, diagrams needed for cutting-edge computations give rise to special functions beyond the ones that live on the Riemann sphere. These functions beyond MPLs are associated to periods of different geometries. The first example being the massive sunrise diagram in even dimensions which can be associated to an elliptic curve (which is isomorphic to a torus) Bloch:2013tra; Broadhurst:1987ei; Bauberger:1994by; Bauberger:1994hx; Laporta:2004rb. Recent progress has been made to go beyond the torus, on the one hand by studying Feynman integrals associated to higher genus Riemann surfaces Georgoudis:2015hca; Duhr:2024uid and their function space DHoker:2023vax; Baune:2024biq; DHoker:2025szl. On the other hand by studying Feynman integrals associated to higher dimensional Calabi-Yau manifolds Primo:2017ipr; Klemm:2019dbm; Bourjaily:2018yfy and their function space Bonisch:2024nru; Duhr:2025ppd; Duhr:2025tdf.
@@
In recent years, substantial effort has been dedicated to extend the definition of canonical differential equations to Feynman integrals beyond polylogarithms, and much progress has been made, in particular in the elliptic and one-parameter Calabi-Yau cases Gorges:2023zgv; Pogel:2022yat; Adams:2016xah; Adams:2018bsn; Pogel:2022ken; Ahmed:2024tsg; Frellesvig:2021hkr; Dlapa:2022wdu; Frellesvig:2023iwr; Chen:2025hzq; Pogel:2022vat. In particular, ref. Pogel:2022vat provides an ansatz to bring equal mass banana integrals, with underlying single-scale Calabi-Yau geometry, at any loop order, into $\epsilon$-factorised form. Moreover, very recently in ref. Duhr:2025lbz, a generalisation of the method of ref. Gorges:2023zgv has been developed for generic integrals with an underlying Calabi-Yau geometry (and beyond), providing explicit calculations for cutting edge Calabi-Yau one-scale integrals. The applicability of this algorithm has been explicitly demonstrated for many state-of-the-art problems containing integrals of elliptic type Gorges:2023zgv; Becchetti:2025rrz; Becchetti:2025oyb; Duhr:2024bzt; Forner:2024ojj; Marzucca:2025eak and for their generalisation to Calabi-Yau varieties Duhr:2024bzt; Forner:2024ojj; Klemm:2024wtd; Driesse:2024feo; Driesse:2024xad; Duhr:2025lbz and higher-genus surfaces Duhr:2024uid.
In this paper, we propose a generalisation of the method introduced in ref. Pogel:2022vat, to multi-scale Calabi-Yau integrals. In the preparation of this work, we have realised that our generalisation coincides with the method from refs. Gorges:2023zgv; Duhr:2025lbz, applied to multi-scale Calabi-Yau Feynman integrals. In fact, this is the first application of refs. Gorges:2023zgv; Duhr:2025lbz to this class of integrals. To support our claim, we derive locally at the MUM-point, for the first time, a canonical differential equation for banana integrals with unequal masses.
This paper is organised as follows. In section 2 we set our conventions and review the instruments that will be needed for the procedure we propose. Following this, in section 3, we present our procedure for multi-variate Feynman integrals with underlying Calabi-Yau geometry. In section 4, we recompute a canonical form for the unequal mass sunrise integral using the method presented before. Finally, in section 5, we apply our construction to compute a canonical form for integrals with underlying multi-variate Calabi-Yau geometry, namely, the three-loop banana with two unequal masses (two-parameters K3) and the four-loop banana with two unequal masses (two-parameters CY three-fold). Additionally, we include three appendices. In the first two we prove the validity of our procedure for the elliptic and Calabi-Yau two-fold case. In the third, we briefly review the modular bootstrap method of Ref. Giroux:2022wav. The main quantitative results will be compiled in the ancillary files sunrise.nb, banana_3_1.nb, banana_2_2.nb and banana_4_1.nb.
@@ 2.1 Feynman integrals and their differential equations
where $\alpha_{jk}\beta_{jk}$ can be chosen to lie in $\{0,\pm 1\}$. Any integral of the form (1) can be reduced to a finite linear combination of master integrals (MIs) which has fixed $\nu_{j}$. Every integral belongs to a sector, defined as the set of integrals from the family that share exactly the same propagators, i.e., $I_{\nu_{1},\dots,\nu_{r}}$ and $I_{\nu^{\prime}_{1},\dots,\nu^{\prime}_{r}}$ belong to the same sector if $\theta(\nu_{i})=\theta(\nu^{\prime}_{i})$, for all $1\leq i\leq r$, where $\theta$ denotes the Heaviside step function, $\theta(x)=1$ if $x>0$ and $\theta(x)=0$ otherwise. Determining these MIs and their associated differential equation can be achieved through the Laporta algorithm, where different software implementations are available Lee:2012cn; Smirnov:2008iw; Smirnov:2019qkx; Klappert:2020nbg; Maierhofer:2017gsa; Peraro:2019svx; Guan:2024byi. This leads to a system of linear first-order differential equations for the vector of MIs $\bf I$ that can be shown to be Kotikov:1991hm; Kotikov:1991pm; Henn:2013pwa; Kotikov:1990kg; Gehrmann:1999as:
@@
${\bf{M}}\left({\bf z};\epsilon\right)$ $=\mathcal{P}\hskip-2.27621pt\exp\left(\epsilon\int_{{\bf z}_{0}}^{{\bf z}}\tilde{\bf A}({\bf z}^{\prime})\right){\bf{M}}({\bf z}_{0};\epsilon)\,$, (6)
@@ 2.2 Feynman integrals and their geometry
In all the examples we consider, we find that the maximal cut only features one geometry. In cases where there are more than one geometry, we expect the differential equation to split into different sub-blocks, provided that we choose a good starting basis as predicted by the MHS. Equivalently as putting the propagators on-shell, a maximal cut of a Feynman integral is obtained by evaluating the integrand in Baikov representation on a contour that encircles all the propagator poles. The integration contour varies according to the geometry of the Feynman integral considered. For instance, at one loop we encounter complex curves of genus zero, namely Riemann spheres. In this case, maximal cuts, in the appropriate integer dimension, evaluate to algebraic functions and the differential one-forms obtained in the differential equation are the so-called dlogdlog-forms, which after integration evaluate to multiple polylogarithms (MPLs) Kummer; Remiddi:1999ew; Goncharov:1995; Goncharov:1998kja; Vollinga:2004sn; Duhr:2011zq.
@@
where $f_{k}(\tau)$ is a modular form of weight $k$ for some congruence subgroup and $g^{(k)}(z,\tau)$ are the coefficients of the expansion of the Kronecker function. A similar analysis applies when the degree of the polynomial under the square root in the maximal cut is three Broedel:2017kkb; Broedel:2017siw. For these elliptic cases, the $W_{1}$ part of the weight filtration coincides with the cohomology $H^{1}(\mathcal{E},\mathbb{Q})$, where $\mathcal{E}$ is the elliptic curve. The complete space of master integrals can be identified with the $W_{2}$ part of the weight filtration. Working modulo linear combinations of the two master integrals in $W_{1}$, it leads us to consider the quotient $W_{2}/W_{1}$, which dimension depends on the number of additional master integrals in the maximal cut, capturing their existence.
@@ 2.3 Picard-Fuchs operators
This equation has $k$ independent solutions and we can obtain a basis using Frobenius method. For every point $z\in\mathbb{C}$, we can construct $k$ linearly independent local solutions. Each of them is given in terms of power series which converge up to the nearest singularity. These local solutions can be analytically continued to multivalued global solutions over the whole parameter space. Unless stated otherwise we will always be looking for the solution around $z_{0}=0$. [^2] For solutions away from zero, we can perform the following substitutions: if $z_{0}\neq\infty$, then $z\rightarrow z^{\prime}=z-z_{0}$ and if $z_{0}=\infty$, then $z\rightarrow z^{\prime}=1/z$ The method starts by solving the indicial equation
@@
which generalises the $\tau$ parameter of an elliptic curve and gives a local (at the MUM-point) canonical variable on the moduli space of Calabi-Yau varieties. The dots denote local analytic terms in $z$. The inverse of this map is also known as the mirror map
@@
$N_{0}=1\,,\quad N_{j+1}=z\dfrac{\partial}{\partial z}\dfrac{1}{N_{j}(\psi_{j}^{(l)})}N_{j}\,$, (23)
@@ 2.4 Hodge filtration & Griffiths transversality
can be evaluated in terms of complete elliptic integral of first kind $K(k)$ and K(1k)K(1-k) for $1\leq z\leq\infty$ through:
@@
Introducing $\tau=\Pi_{12}(z)/\Pi_{11}(z)$ and using it as the complex modulus, we can express the periods of the elliptic curve as modular forms of weight one in $\tau$. Hence, we expect at least two master integrals in the maximal cut of a Feynman integral associated with an elliptic curve. This also matches with the considerations from section 2.2.
@@
Hence, for the Calabi-Yau manifolds that have appeared in the Feynman integrals literature, to span the middle cohomology $\mathcal{H}^{l}(M_{l},\mathbb{C})$ one can just take a basis of derivatives of the holomorphic $l$-form $\Omega$. [^5] Note that this is not always the case: for Calabi-Yau $l$-folds with $l\geq 4$, taking derivatives of the holomorphic form is not always enough to span the full middle cohomology, i.e., manifolds with Hodge numbers (1, 1,h, 1, 1){(1,\,1,\,h,\,1,\,1)} with $h>1$. The examples we consider are at most Calabi-Yau three-folds so we do not encounter this problem. Note that all differentials but $\Omega$ are meromorphic. This will prove very useful when looking for a good starting basis for Feynman integrals with Calabi-Yau geometry, as the $W_{l}$ part of the weight filtration coincides with $H^{l}(\mathrm{CY},\mathbb{Q})$. This implies that the maximal cut will contain at least $n=\mathrm{dim}(H^{l}(\mathrm{CY},\mathbb{Q}))$ master integrals, that correspond to the holomorphic period and its independent derivatives for $\epsilon\to 0$. As an example, let us return to the elliptic curves from Eq. (32). The derivative with respect to $z$ of $\Omega=\mathrm{d}x/y$, can be written, up to exact terms, as a linear combination of $\mathrm{d}x/y$ and $x\,\mathrm{d}x/y$, where the latter is a differential of second kind. Pairing this differential of second kind with the cycles (33) will yield the complete elliptic integrals of second kind $E(k^{2})$ and $E(1-k^{2})$ for $1\leq z\leq\infty$ through:
@@
$\mathbf{\Pi}(\mathbf{z})^{T}\mathbf{\Sigma}\partial_{\bf z}^{\bf k}\mathbf{\Pi}(\mathbf{z})=\int_{M_{l}}\Omega\,\wedge\,\partial_{\bf z}^{\bf k}\Omega(\mathbf{z})=\begin{cases}0&\text{for }0\leq|k|<l,\\ C_{\mathbf{k}}(\mathbf{z})&\text{for }|k|=l,\end{cases}$ (42)
@@
which implies that we can find $2\hat{h}+2$ independent periods
@@
where the subscript determines the logarithmic power of the period in the Frobenius basis. These periods have been chosen by requiring the intersection matrix $\Sigma$ to be antidiagonal [^7] Where here, the identity matrices here should be understood as antidiagonal, e.g. $\mathds{1}_{2\times 2}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}\,$.
$\mathbf{\Sigma}=\begin{pmatrix}0&\mathds{1}_{(\hat{h}+1)\times(\hat{h}+1)}\\ -\mathds{1}_{(\hat{h}+1)\times(\hat{h}+1)}&0\end{pmatrix}\,$. (47)
@@ 2.5 Abelian differentials of first, second and third kind
where we pick the root ordering $1<z<\infty$. Topologically this corresponds to two Riemann spheres connected by two branch cuts, which we can choose to go from 0 to $1$ and from $z$ to $\infty$. If we allow differentials of the third kind, with non-vanishing residue at a marked point $c$, this corresponds to a puncture on each of the Riemann spheres at the marked point $c$. Explicitly one gets the following forms of first, second and third kind:
[…]
@@ 2.6 $\epsilon$-factorising L-loop banana integrals of equal mass
As we have reviewed in section 2.4, a good basis for master integrals in $D=2$ is the derivative basis, that for a $\text{L}$-loop Banana integral takes the form:
@@ 3 Algorithm to construct a canonical basis
First of all, we need to identify the geometry of the block we are looking at. This can either be done with a Baikov analysis, or finding the periods annihilated by the Picard-Fuchs operator. For cases like the banana integrals in two dimensions, the corner integral ($I_{1,\dots,1}$) evaluates to periods of a Calabi-Yau manifold on its maximal cut in $D=2$ dimensions. This is, however, not always the case and one may have to shift dimensions or take linear combinations of the available master integrals to make the underlying geometry manifest.
@@
where $z_{i}$ are variables from which we can write a Calabi-Yau PF operator for $I_{1}$ which has a MUM-point at $z_{i}\rightarrow 0$. [^9] For elliptics we have only one $z_{i}$ because there is only one canonical variable $\tau$. Note in particular that there are no derivative terms $\partial_{z_{i}}^{k}I_{1}$ with $k\geq 2$. For Feynman integrals with an underlying elliptic curve, this corresponds to a non-vanishing residue around a simple pole that is not one of the roots of the zero locus of the maximal-cut geometry as described in section 2.5. For instance, Eq. (65) has two master integrals corresponding to the first and second kind differentials which saturate the middle cohomology as evidenced by the Hodge diamond in Eq. (30), and the last master integral must be a differential of the third kind, since we want it to decouple. However, even in elliptic examples, see, for example, section 4.4 in ref. Gorges:2023zgv, not all the decoupling integrals have been chosen such that their maximal cut is an integral of a differential form of the third kind when $\epsilon=0$. Nevertheless, in that case, the differential equation takes the form of Eq. (81). Hence, when the integrand analysis is too involved, or finding independent candidates of the third kind is too complicated, we choose master integrals whose differential equation satisfy Eq. (81). However, this is not the most generic form, for example in section 7.2.2 in ref. Duhr:2025lbz, the decoupling master integral satisfies a more generic differential equation. This is because, unlike integrals of the third kind, this integral has a double pole. In that case, we see that we need more functions than the ones that we use in our examples. Therefore the bases proposed in this paper is valid for decoupling integrals whose differential equation takes the form of Eq. (81). If such a candidate cannot be found, we follow the indications given in ref. Gorges:2023zgv; Duhr:2025lbz.
@@
From the examples we have computed, we have observed that the basis we propose has the following properties:
@@
Hence, we call it canonical. [^10] So far there is no clear consensus on a definition of canonical differential equation beyond MPLs. In ref. Duhr:2025lbz, a definition was proposed, which includes the independence of the forms in addition to our defining properties. Checking the independence of the forms goes beyond the scope of this paper.
@@
$\mathbf{I}=(\underbrace{I_{1},I_{2},\dots,I_{n-1},I_{n}}_{\text{first, second kind integrals}},\underbrace{I^{\text{dec}}_{n+1},\dots,I^{\text{dec}}_{n+m}}_{\text{decoupling integrals}})\,$, (82)
@@
and we additionally require the differential equation for $\mathbf{I}_{\mathbf{B}}$ to take the form in Eq. (81). Now let us concretely show how this works for $l=1,2,3$ and then propose a generalisation for l>3{l>3}.